OpenStudy (chris215):

Write the equation of the function g(x) if g(x) = f(x + 2) + 4 and f(x) = x3 - 7.

OpenStudy (jango_in_dtown):

What is x3?

OpenStudy (jango_in_dtown):

@chris215

OpenStudy (chris215):

x^3

OpenStudy (jango_in_dtown):

ok. so we have g(x)=f(x + 2) + 4 =(x+2)^3-7+4=(x+2)^3-3

OpenStudy (chris215):

thank you!